# finite free probability and $S$ transforms of Jacobi processes

### v1

@@ 3.3. Analysis of the Characteristic curves for $\lambda=1,\theta=1/2$
In this paragraph, we restrict our attention to the special parameter set $\theta=1/2$, $\lambda=1$, and assume for sake of simplicity that $\mathcal{S}^{(1,1/2)}_{0}(z)=1$ (in which case the initial spectral distribution of the free Jacobi process is $\delta_{1}$). This restriction is mainly motivated by the fact that the moment-generating function $\mathcal{M}_{t}^{(1,1/2)}$ admits an explicit expression, as proved in [DHH12], from which a series expansion of $\mathcal{S}_{t}^{(1,1/2)}$ was derived in [Dem17] making use of Lagrange inversion formula. As we shall see below, the analysis of the Characteristic curves of the pde (4) in the special case corresponding to $\lambda=1,\theta=1/2$ leads to a considerably more tractable expression of $\mathcal{S}^{(1,1/2)}_{t}$ than the one displayed in Proposition 5.2. from [Dem17].
@@
Let $t\geq 0$. The S-transform $\mathcal{S}_{t}^{(1,1/2)}$ is given locally around $z=0$ by
@@
For any time $t>0$, $\xi_{2t}$ is the inverse, in a vicinity of $u=1$, of the Herglotz transform of the spectral distribution of the free unitary Brownian motion at time $2t$ [Bia97]:
@@
while(7) yields again:
@@
Now, recall from [Bia97], Lemma 12, that for any fixed time $t$, there exists a Jordan $\Gamma_{t}$ lying in the right half-plane and containing $w=1$ where $\xi_{t}$ is a one-to-one map onto the open unit disc. As a matter of fact, there exists a neighbourhood of the origin such that for any $z$ there, the associated characteristic curve $s\mapsto z(s)$ is globally defined in time.
@@ 4. Averaged characteristic polynomial of the Hermitian Jacobi Process
  • In Proposition (4.1.1), we revisit Voit’s result showing that the averaged characteristic polynomial satisfies the inverse Jacobi-heat equation for $p\wedge q>m-(1/2)$. In doing so, we prove it directly using stochastic calculus, without relying on elementary symmetric polynomials. We also emphasise that this result holds for the larger set $p\wedge q>m-1$ by virtue of Corollary 7 in [GM14].
@@
  • Proposition 4.2.1 is the finite free analogue of Corollary $3.$ proved in [DHH12]: when $\lambda=1,\theta=1/2$, the image of the roots of the averaged characteristic polynomial of the Hermitian Jacobi process by the transformation $x\mapsto 2\arccos(\sqrt{x})$ coincides with the roots of the Hermite unitary polynomial.
  • Proposition 4.3.1 provides an expansion of $\chi_{t}^{(r,s,m)}$ which is valid for any initial value of the eigenvalues process (1). While Proposition (4.1.1) also shows that such expansion follows from the inverse Jacobi-heat equation, the one provided by Proposition 4.3.1 relies on the so-called dual Cauchy identity satisfied by multivariate Jacobi polynomials. This identity originates in algebraic combinatorics (see [OO12] and references therein) and opens the way to use the mutual orthogonality of the multivariate Jacobi polynomials.
@@ 4.1. Averaged Characteristic Polynomial
$F(x;u_{1},\dots,u_{m})=\prod_{i=1}^{m}(x-u_{i}),\quad x\neq u_{i}\in[0,1],$
then the averaged characteristic polynomial of $X_{t}$ is given by the integral:
@@
In the series of papers [Voi20], [Voi23], [Voi25], the roots of the characteristic polynomial of the (rescaled) $\beta$-Jacobi particle system valued in [1,1][-1,1] are studied, among others. In particular, Corollary 3.4 from [Voi20] shows that the averaged elementary symmetric polynomials in this random system do not depend on $\beta$. Besides, Theorem 4.4 from [Voi25] describes the dynamics of the roots of the averaged characteristic polynomial through an inverse Jacobi-heat equation. Though the $\beta$-Jacobi particle systems in $[0,1]$ and in [1,1][-1,1] are related by an affine transformation, the relation between elementary symmetric polynomials in both corresponding roots is far from being trivial. Nonetheless, one expects that the dynamics of the roots associated with the Jacobi particle system in $[0,1]$ are still governed by an inverse Jacobi-heat equation. In this respect, a direct application of Itô’s formula together with appropriate algebraic transformations shows that this is indeed true.
@@
$\partial_{t}\chi_{t}^{(r,s,m)}(x)=-\left\{\mathcal{L}_{x}^{(r,s)}+m(r+s+m+1)\right\}\chi_{t}^{(r,s,m)}(x).$
@@
It suffices to find the expansion of $(x-1)^{m}$ in the Jacobi polynomial basis. But this is afforded by the following formula [KK99]:
@@
$(1-x)^{m}=m!\sum_{j=0}^{m}\frac{\Gamma(s+m+1)\Gamma(r+s+j+1)(r+s+2j+1)}{(m-j)!\Gamma(j+s+1)\Gamma(r+s+m+j+2)}Q_{j}^{(r,s)}(x)$
@@
Applying the shifted Jacobi-heat semi-group:
$e^{-t(\mathcal{L}_{x}^{(r,s)}+m(d-m+1))},$
to the obtained (finite) expansion, we are done. ∎
@@
is proportional to the monic Jacobi polynomial $Q_{m}^{(r,s)}(x)/k_{m}^{(r,s)}$, where $k_{m}^{(r,s)}$ is the leading term of $Q_{m}^{(r,s)}(x)$. This is in agreement with the fact that the averaged characteristic polynomial of the Jacobi unitary ensemble, the weak limit of the Hermitian Jacobi process as $t\rightarrow+\infty$, is given by a Jacobi polynomial. Indeed, the eigenvalues of the JUE (Jacobi Unitary Ensemble) are given by a Selberg weight, as such its averaged characteristic polynomial is an instance of the celebrated Aomoto integral (see e.g. [LT03]).
@@
The first part of this proposition is due to Voit, as stated in Theorem 4.4 of [Voi25]. Now, let’s turn to the second statement. We have
@@
since $\lambda_{k}(t)>0$ almost surely for any $1\leq k\leq m$ [Dem10, GM14]. It follows that $x_{k}^{(r,s,m)}(t)\neq 0$ for any $1\leq k\leq m$. Similarly,
@@ 4.2. Relation to the Hermite unitary polynomial
In [DHH12], it was shown that the spectral distribution of the free Jacobi process with parameters $(1,1/2)$ in the compressed probability space coincides with the spectral distribution of
@@
where $(Y_{t})_{t\geq 0}$ is the free unitary Brownian motion [Bia97]. Note in this respect that this special set of parameters corresponds at the matrix level to Hermitian Jacobi processes, which are radial parts of ‘asymptotically square’ (as $m\rightarrow+\infty$) principal minors of a unitary Brownian motion whose size is asymptotically twice the sizes of these corners.
On the other hand, the finite analogue of the free unitary convolution, with respect to which $(Y_{t})_{t\geq 0}$ is a free Lévy process, was introduced and studied in [Mir21]. This finite convolution is encoded by the zeroes of the Hermite Unitary Polynomial defined for any time $t\geq 0$ by [Kab25] [^2] This definition differs from the original one given in [Mir21] by the time change $t\mapsto-t(n-1)$.:
@@
In particular, Lemma 2.1. in [Kab25] asserts that the zeroes of $H_{d}(\cdot,t)$ lie on the unit circle and we infer from Corollary 3.35 proved in [Mir21] that the dynamics of the corresponding angles satisfy the following ODE:
@@
In particular, Theorem 6.1. in [Voi25] shows that these angles (and in turn the roots of $H_{d}(\cdot,t)$) remain distinct for all times $t>0$ even if they collapse at $t=0$. Besides, since $H_{d}(\cdot,t)$ has real coefficients, then its roots are pairwise-conjugate. Note also that z=1z=-1 is not a root of $H_{d}(\cdot,t)$ since $H_{d}(-1,t)$ is a sum of positive terms, while $z=1$ is so only when the degree $d$ is odd. Based on this discussion, it is tempting to wonder whether both root dynamics (12) and (11) are related through the transformation [^3] For sake of simplicity, we omit the dependence on the parameters.:
@@
This is the ODE (11) with p=q=m(1/2)p=q=m-(1/2). But Theorem 1.1. in [AVW24] shows (after performing an affine variable change) that (11) admits a unique solution for any p,q>m1p,q>m-1 which remains in the open domain $\{0<x_{m}<\cdots<x_{1}<1\}$ at any time $t>0$. It follows that
@@ 4.3. A more general expansion
We derive a more general expansion of the averaged characteristic polynomial $\chi_{t}^{(r,s,m)}$ which is valid for any initial value $w=\lambda_{0}$ of the eigenvalues process. Our main ingredients are the heat kernel of the latter and the dual Cauchy identity satisfied by the (symmetric) multivariate Jacobi polynomials recalled below. These polynomials are mutually orthogonal with respect to the unitary Selberg weight. This property is not satisfied by any orthogonal set of multivariable polynomials since it requires the orthogonality of any two polynomials corresponding to different partitions. For the sake of completeness, we provide a brief reminder of the key facts and results necessary to prove the expansion below. We refer the reader to the paper [Dem10] for more details.
@@
Recall from [OO12] the dual Cauchy-identity:
@@
Proposition 7.1 in [OO12] gives an explicit expression of $Q_{\tau}^{(r,s,m)}(1^{m})$ as a ratio of products of Gamma functions. After lengthy (but easy) computations, one retrieves the second statement of Corollary 4.1.2. Moreover, if $(w_{i}=z_{i}^{(r,s)})_{1\leq i\leq m}$ are the zeroes of the Jacobi polynomials $Q_{m}^{(r,s)}$, then $Q_{(1^{m-j})}^{(r,s,m)}(w)=0$ for all $0\leq j\leq m-1$ since $Q_{\tau_{1+m-1}}^{(r,s)}(z_{i})=0$ for any $1\leq i\leq m$. Consequently, $\chi_{t}^{(r,s,m)}(x)=Q_{m}^{(r,s)}(x)/k_{m}^{(r,s)}$ for any $t\geq 0$ and agrees with the fact that $(z_{i})_{1\leq i\leq m}$ is the stationary solution of (11) (see Proposition 7 in [Voi23]).
@@ 5. Frozen Hermitian Jacobi process and finite free probability
In this section, we study the Frozen Jacobi process and the finite S transform. The main results are
  • The convergence of the counting measure of the roots of $\chi_{t}^{(r,s,m)}$ to the distribution of the free Jacobi process at time $t\geq 0$. Stated in Corollary 5.1.2, it follows from the tightness of this counting measure together with the differential system derived in Proposition 5.1.1 and satisfied by its moment sequence.
  • Theorem 5.2.1 where we establish a general result, of independent interest, concerning the convergence of the finite differences of the finite free $S$ transform.
  • Proposition 5.3.2 and Theorem 5.3.3 where the PDE in Theorem 3.1.1 is derived as a limit of an “ODE with finite differences” for the finite free $T$ transform.
@@ 5.1. High-dimensional regime
$m_{\ell}^{(r,s,m)}(t)=\mu^{(r,s,m)}_{t}(x^{\ell}),\quad\ell\geq 0,$
be the moment sequence of $\mu_{t}^{(r,s,m)}$, then:
@@
The moments $(m_{\ell}^{(r,s,m)}(t))_{\ell\geq 0}$ are the solutions of the following differential system :
@@
Appealing to the ODE (11), we readily derive
$\frac{d}{dt}m_{\ell}^{(r,s,m)}$ $=\frac{\ell}{m}\sum_{i=1}^{m}(x_{t}^{i})^{\ell-1}\frac{d}{dt}x_{t}^{i}$
$=\frac{\ell}{m}\left(mpm_{\ell-1}^{(r,s,m)}-m(p+q)m_{\ell}^{(r,s,m)}+\sum_{i=1}^{m}\sum_{j\neq i}\frac{(x_{t}^{i})^{\ell}(1-x_{t}^{j})+(x_{t}^{i})^{\ell-1}x_{t}^{j}(1-x_{t}^{i})}{x^{i}_{t}-x_{t}^{j}}\right).$
@@
$\sum_{i=1}^{m}\sum_{j\neq i}\frac{(x_{t}^{i})^{\ell}(1-x_{t}^{j})+(x_{t}^{i})^{\ell-1}x_{t}^{j}(1-x_{t}^{i})}{x^{i}_{t}-x_{t}^{j}}=\sum_{i=1}^{m}\sum_{j\neq i}\frac{-2(x_{t}^{i})^{\ell}x_{t}^{j}}{x^{i}_{t}-x_{t}^{j}}+\frac{(x_{t}^{i})^{\ell}+(x_{t}^{i})^{\ell-1}x_{t}^{j}}{x^{i}_{t}-x_{t}^{j}},$
@@
We now proceed to the aforementioned convergence result. To this end, recall from [Dem08] that the moments
@@
$\tilde{m}_{\ell}^{(r,s,m)}(t):=m_{\ell}^{(r,s,m)}\left(\frac{t}{d}\right)$
converges to $m_{\ell}(t)$ as well (in the same regime).
@@
Since $\ell$ is fixed, then the right-hand side of (17) is uniformly bounded for sufficiently large $m,p(m),d(m)$. Consequently, $\|\partial_{t}\tilde{m}_{\ell}^{(r(m),s(m),m)}\|_{\infty}$ is so whence equicontinuity follows. Hence $\tilde{m}^{(r(m),s(m),m)}$ converges locally uniformly in $t$ and in turn so does $d\tilde{m}^{(r(m),s(m),m)}/dt$. Passing to the limit in (17) and using induction on $\ell$, we conclude that any limiting moment sequence satisfies (16) for any $\ell$. Since the latter admits a unique solution for any fixed intial data at $t=0$, there is one and only one accumulation point for $\tilde{m}_{\ell}^{(r(m),s(m),m)}(t)$ with prescribed initial data. The proposition is proved. ∎

### v3

@@ 3.3. Analysis of the Characteristic curves for $\lambda=1,\theta=1/2$
In this paragraph, we restrict our attention to the special parameter set $\theta=1/2$, $\lambda=1$, and assume for sake of simplicity that $\mathcal{S}^{(1,1/2)}_{0}(z)=1$ (in which case the initial spectral distribution of the free Jacobi process is $\delta_{1}$). This restriction is mainly motivated by the fact that the moment-generating function $\mathcal{M}_{t}^{(1,1/2)}$ admits an explicit expression, as proved in [MR3071702], from which a series expansion of $\mathcal{S}_{t}^{(1,1/2)}$ was derived in [MR3679612] making use of Lagrange inversion formula. As we shall see below, the analysis of the Characteristic curves of the pde (4) in the special case corresponding to $\lambda=1,\theta=1/2$ leads to a considerably more tractable expression of $\mathcal{S}^{(1,1/2)}_{t}$ than the one displayed in Proposition 5.2. from [MR3679612].
@@
Let $t\geq 0$. The S transform $\mathcal{S}_{t}^{(1,1/2)}$ is given locally around $z=0$ by
@@
For any time $t>0$, $\xi_{2t}$ is the inverse, in a vicinity of $u=1$, of the Herglotz transform of the spectral distribution of the free unitary Brownian motion at time $2t$ [MR1430721]:
@@
while equation (7) yields again:
@@
Now, recall from [MR1430721], Lemma 12, that for any fixed time $t$, there exists a Jordan $\Gamma_{t}$ lying in the right half-plane and containing $w=1$ where $\xi_{t}$ is a one-to-one map onto the open unit disc. As a matter of fact, there exists a neighbourhood of the origin such that for any $z$ there, the associated characteristic curve $s\mapsto z(s)$ is globally defined in time.
@@ 4. Averaged characteristic polynomial of the Hermitian Jacobi Process
  • In Proposition (4.1.1), we revisit Voit’s result showing that the averaged characteristic polynomial satisfies the inverse Jacobi-heat equation for $p\wedge q>m-(1/2)$. In doing so, we prove it directly using stochastic calculus, without relying on elementary symmetric polynomials. We also emphasise that this result holds for the larger set $p\wedge q>m-1$ by virtue of Corollary 7 in [MR3296535].
@@
  • Proposition 4.2.1 is the finite free analogue of Corollary 3.3. proved in [MR3071702]: when $\lambda=1,\theta=1/2$, the image of the roots of the averaged characteristic polynomial of the Hermitian Jacobi process by the transformation $x\mapsto 2\arccos(\sqrt{x})$ coincides with the roots of the Hermite unitary polynomial.
  • Proposition 4.3.1 provides an expansion of $\chi_{t}^{(r,s,m)}$ which is valid for any initial value of the eigenvalues process (1). While Proposition (4.1.1) also shows that such expansion follows from the inverse Jacobi-heat equation, the one provided by Proposition 4.3.1 relies on the so-called dual Cauchy identity satisfied by multivariate Jacobi polynomials. This identity originates in algebraic combinatorics (see [MR3075094] and references therein) and opens the way to use the mutual orthogonality of the multivariate Jacobi polynomials.
@@ 4.1. Averaged Characteristic Polynomial
$F(x;u_{1},\dots,u_{m})=\prod_{i=1}^{m}(x-u_{i}),\quad x\in[0,1],$
then the averaged characteristic polynomial of $X_{t}$ is:
@@
In the series of papers [MR4092353], [MR4673381], [MR4781073], the roots of the characteristic polynomial of the (rescaled) $\beta$-Jacobi particle system valued in [1,1][-1,1] are studied, among others. In particular, Corollary 3.4 from [MR4092353] shows that the averaged elementary symmetric polynomials in this random system do not depend on $\beta$. Besides, Theorem 4.4 from [MR4781073] describes the dynamics of the roots of the averaged characteristic polynomial through an inverse Jacobi-heat equation. Though the $\beta$-Jacobi particle systems in $[0,1]$ and in [1,1][-1,1] are related by an affine transformation, the relation between elementary symmetric polynomials in both corresponding roots is far from being trivial. Nonetheless, one expects that the dynamics of the roots associated with the Jacobi particle system in $[0,1]$ are still governed by an inverse Jacobi-heat equation. In this regard, a direct application of Itô’s formula together with appropriate algebraic transformations shows that this is indeed true.
@@
$\partial_{t}\chi_{t}^{(r,s,m)}(x)=-\left\{\mathcal{L}_{x}^{(r,s)}+m(r+s+m+1)\right\}\chi_{t}^{(r,s,m)}(x),$
@@
It suffices to find the expansion of $(x-1)^{m}$ in the Jacobi polynomial basis. But this is afforded by the following formula [MR1697414]:
@@
$(1-x)^{m}=m!\sum_{j=0}^{m}\frac{\Gamma(s+m+1)\Gamma(r+s+j+1)(r+s+2j+1)}{(m-j)!\Gamma(j+s+1)\Gamma(r+s+m+j+2)}Q_{j}^{(r,s)}(x)$ (10)
@@
Applying the shifted Jacobi-heat semi-group $e^{-t(\mathcal{L}_{x}^{(r,s)}+m(d-m+1))}$ to (10), we are done. ∎
@@
is proportional to the monic Jacobi polynomial $Q_{m}^{(r,s)}(x)/k_{m}^{(r,s)}$, where $k_{m}^{(r,s)}$ is the leading term of $Q_{m}^{(r,s)}(x)$. This is in agreement with the fact that the averaged characteristic polynomial of the Jacobi unitary ensemble, the weak limit of the Hermitian Jacobi process as $t\rightarrow+\infty$, is given by a Jacobi polynomial. Indeed, the eigenvalues of the JUE (Jacobi Unitary Ensemble) are given by a Selberg weight, as such its averaged characteristic polynomial is an instance of the celebrated Aomoto integral (see e.g. [MR1985318]).
@@
The first part of this proposition is due to Voit, as stated in Theorem 4.4 of [MR4781073]. Now, let’s turn to the second statement. We have
@@
since $\lambda_{k}(t)>0$ almost surely for any $1\leq k\leq m$ [MR2719370, MR3296535]. It follows that $x_{k}^{(r,s,m)}(t)\neq 0$ for any $1\leq k\leq m$. Similarly,
@@ 4.2. Relation to the Hermite unitary polynomial
In [MR3071702], it was shown that the spectral distribution of the free Jacobi process with parameters $(1,1/2)$ in the compressed probability space coincides with the spectral distribution of
@@
where $(Y_{t})_{t\geq 0}$ is the free unitary Brownian motion [MR1430721]. Note in this respect that this special set of parameters corresponds at the matrix level to Hermitian Jacobi processes, which are radial parts of ‘asymptotically square’ (as $m\rightarrow+\infty$) principal minors of a unitary Brownian motion whose size is asymptotically twice the sizes of these truncations.
On the other hand, the finite analogue of the free unitary convolution, with respect to which $(Y_{t})_{t\geq 0}$ is a free Lévy process, was introduced and studied in [MR4285332]. This finite convolution is encoded by the zeroes of the Hermite unitary polynomial defined for any time $t\geq 0$ by [MR4912666] [^2] This definition differs from the original one given in [MR4285332] by the time change $t\mapsto-t(n-1)$.:
@@
In particular, Lemma 2.1. in [MR4912666] asserts that the zeroes of $H_{d}(\cdot,t)$ lie on the unit circle and we infer from Corollary 3.35 proved in [MR4285332] that the dynamics of the corresponding angles satisfy the following ODE:
@@
In particular, Theorem 6.1. in [MR4781073] shows that these angles (and in turn the roots of $H_{d}(\cdot,t)$) remain distinct for all times $t>0$ even if they collapse at $t=0$. Besides, since $H_{d}(\cdot,t)$ has real coefficients, then its roots are pairwise-conjugate. Note also that z=1z=-1 is not a root of $H_{d}(\cdot,t)$ since $H_{d}(-1,t)$ is a sum of positive terms, while $z=1$ is so only when the degree $d$ is odd. Based on this discussion, it is tempting to wonder whether both root dynamics (13) and (12) are related through the transformation [^3] For sake of simplicity, we omit the dependence on the parameters.:
@@
This is the ODE (12) with p=q=m(1/2)p=q=m-(1/2). But Theorem 1.1. in [MR4751304] shows (after performing an affine variable change) that (12) admits a unique solution for any p,q>m1p,q>m-1 which remains in the open domain $\{0<x_{m}<\cdots<x_{1}<1\}$ at any time $t>0$. It follows that
@@ 4.3. A more general expansion
We derive a more general expansion of the averaged characteristic polynomial $\chi_{t}^{(r,s,m)}$ which is valid for any initial value $w=\lambda_{0}$ of the eigenvalues process. Our main ingredients are the heat kernel of the latter and the dual Cauchy identity satisfied by the (symmetric) multivariate Jacobi polynomials recalled below. These polynomials are mutually orthogonal with respect to the unitary Selberg weight. This property is not satisfied by any orthogonal set of multivariable polynomials since it requires the orthogonality of any two polynomials corresponding to different partitions. For the sake of completeness, we provide a brief reminder of the key facts and results necessary to prove the expansion below. We refer the reader to the paper [MR2719370] for more details.
@@
Recall from [MR3075094] the dual Cauchy-identity:
@@
Proposition 7.1 in [MR3075094] gives an explicit expression of $Q_{\tau}^{(r,s,m)}(1^{m})$ as a ratio of products of Gamma functions. After lengthy (but easy) computations, one retrieves the second statement of Corollary 4.1.2. Moreover, if $(w_{i}=z_{i}^{(r,s)})_{1\leq i\leq m}$ are the zeroes of the Jacobi polynomials $Q_{m}^{(r,s)}$, then $Q_{(1^{m-j})}^{(r,s,m)}(w)=0$ for all $0\leq j\leq m-1$ since $Q_{\tau_{1+m-1}}^{(r,s)}(z_{i})=0$ for any $1\leq i\leq m$. Consequently, $\chi_{t}^{(r,s,m)}(x)=Q_{m}^{(r,s)}(x)/k_{m}^{(r,s)}$ for any $t\geq 0$ and agrees with the fact that $(z_{i})_{1\leq i\leq m}$ is the stationary solution of (12) (see Proposition 7 in [MR4673381]).
@@ 5. Frozen Hermitian Jacobi process and finite free probability
In this section, we study the Frozen Jacobi process and its finite S transform. The main results are
  • The convergence of the counting measure of the roots of $\chi_{t}^{(r,s,m)}$ to the distribution of the free Jacobi process at time $t\geq 0$. Stated in Corollary 5.1.2, it follows from the differential system derived in Proposition 5.1.1 and satisfied by the moment sequence of this counting measure, together with their relative compactness.
  • Theorem 5.2.1 where we establish a general result, of independent interest, regarding the convergence of the finite differences of the finite free $S$ transform.
  • Corollary 5.3.2 and Theorem 5.3.3 where the PDE in Theorem 3.1.1 is derived as a limit of an “ODE with finite differences” for the finite free $T$ transform.
@@ 5.1. High-dimensional regime
$m_{\ell}^{(r,s,m)}(t)(t)=\mu^{(r,s,m)}_{t}(x^{\ell}),\quad\ell\geq 0,$
be the moment sequence of $\mu_{t}^{(r,s,m)}$.
@@
For any $r,s$ and $m\geq 1$ an integer, the trajectory of the infinite sequence of moments $(m_{\ell}^{(r,s,m)}(t)(t))_{\ell\geq 0}$ is the solution of the following lower triangular differential system :
@@
Appealing to the ODE (12), we derive
$\frac{d}{dt}m_{\ell}^{(r,s,m)}(t)(t)$ $=\frac{\ell}{m}\sum_{i=1}^{m}(x^{i}(t))^{\ell-1}\frac{d}{dt}x^{i}(t)$
$=\frac{\ell}{m}\bigg(mpm_{\ell-1}^{(r,s,m)}(t)-m(p+q)m_{\ell}^{(r,s,m)}(t)(t)$
$\hskip 56.9055pt+\sum_{i=1}^{m}\sum_{j\neq i}\frac{(x^{i}(t))^{\ell}(1-x^{j}(t))+(x^{i}(t))^{\ell-1}x^{j}(t)(1-x^{i}(t))}{x^{i}(t)-x^{j}(t)}\bigg).$
@@
$\sum_{i=1}^{m}\sum_{j\neq i}\frac{(x^{i}(t))^{\ell}(1-x^{j}(t))+(x^{i}(t))^{\ell-1}x^{j}(t)(1-x^{i}(t))}{x^{i}(t)-x^{j}(t)}$
$\hskip 142.26378pt=\sum_{i=1}^{m}\sum_{j\neq i}\frac{-2(x^{i}(t))^{\ell}x^{j}(t)}{x^{i}(t)-x^{j}(t)}+\frac{(x^{i}(t))^{\ell}+(x^{i}(t))^{\ell-1}x^{j}(t)}{x^{i}(t)-x^{j}(t)},$
@@
We now proceed to the aforementioned convergence result. To this end, recall from [MR2384475] that the moments
@@
$\tilde{m}_{\ell}^{(r(m),s(m),m)}(t):=m_{\ell}^{(r(m),s(m),m)}(t)\left(\frac{t}{d(m)}\right)$
converges to $m_{\ell}(t)$ as well.
@@
Since $\ell$ is fixed, then the right-hand side of (18) is uniformly bounded for sufficiently large $m,p(m),d(m)$. Consequently, $\|\partial_{t}\tilde{m}_{\ell}^{(r(m),s(m),m)}\|_{\infty}$ is so whence equicontinuity follows. By Arzelà-Ascoli Theorem, the moments sequence $\{\tilde{m}^{(r(m),s(m),m)}\}_{m}$ converges locally uniformly in $t$ along a subsequence and in turn so does $d\tilde{m}^{(r(m),s(m),m)}/dt$. Passing to the limit in (18) and using induction on $\ell$, we conclude that any limiting moment sequence satisfies (17) for any $\ell$. Since the latter admits a unique solution for any fixed intial data at $t=0$, there is one and only one accumulation point for $\tilde{m}_{\ell}^{(r(m),s(m),m)}(t)$ with prescribed initial data. The proposition is proved. ∎